OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 256 and a3 = 16?

OpenStudy (anonymous):

\[a_1=256\] \[a_3=16\] \[a_3/a_1=r^2\] \[16/256=1/16=r^2\] \[a_7=a_1*r^6\] \[a_7=256*(1/16)^3=1/16\] \[1/16\]

OpenStudy (anonymous):

So 0.0625...thankyou sooo much!!